# From Simple to Complex: Knowledge Transfer in Safe and Efficient Reinforcement Learning for Autonomous Driving

### v1

@@ 3.3.2 High-level Decision-Making Layer
Table 1: Parameters of S2CD framework
Hyper-parameters Value Hyper-parameters Value
Maximal learning rate 0.0005 Discount factor $\gamma$ 0.96
Learning rate decay True Lambda entropy $\beta$ 0.01
Total steps of one episode 5,000 Clip parameter $\epsilon$ 0.2
Total training steps 300K Lambda advantage $\lambda$ 0.98
Optimizer AdamW Hyperparameter $q_{1}$ 3
Mini batch size 64 Hyperparameter $q_{2}$ 10
Hyperparameter of AC $\psi$ 0.2 Initial Lagrange multiplier $\xi$ 0.01
Efficiency weight $\alpha_{1}$ 0.5 Safety weight $\alpha_{2}$ 1.0
Tolerance coefficient $\varepsilon$ 0.5
@@ 5.1 Model Training
Figure 4: The training curves for all algorithms are presented. Each algorithm is trained with 3 different random seeds, accumulating a total of 500K training steps for each case, except for S2CD, which requires only 300K training steps. To ensure a comprehensive assessment of algorithm performance, 2 evaluation episodes are conducted for every 5,000 training steps. The average value of these 2 evaluation episodes is taken as the result.
@@ 5.2 Model Evaluation
Figure 5: The curves of model evaluation for all algorithms. Each model trained with varying random seeds is evaluated twice, and each algorithm undergoes 6 evaluation runs, with each evaluation consisting of 200 episodes.
@@ 5.3.1 The Impact of Different Teacher Quality
Figure 6: The training and evaluation processes involve different teacher models. 1. Model Training: Each algorithm is trained with 3 different random seeds, accumulating a total of 300K training steps for each case. Two evaluation episodes are conducted for every 5,000 training steps. The average value of these 2 evaluation episodes is taken as the result. During model training, we record the return value, safety cost, and collision counts. 2. Model Evaluation: Each model trained with varying random seeds is evaluated twice. Each algorithm undergoes 6 evaluation runs, with each evaluation consisting of 200 episodes. For model evaluation, we record the success rate.
Table 10: Total rewards, average speed, and success rate for different $\omega$
@@ 5.3.2 Ablations of Guardian Mechanism
Figure 7: The training and evaluation processes of the ablation experiment. 1. Model Training: Each algorithm is trained with 3 different random seeds, accumulating a total of 300K training steps for each case. Two evaluation episodes are conducted for every 5,000 training steps, and the average value of these 2 episodes is taken as the result. During model training, we record the return value, safety cost, speed, and collision counts. 2. Model Evaluation: Each model trained with varying random seeds is evaluated twice, and each algorithm undergoes 6 evaluation runs, with each evaluation consisting of 200 episodes. For model evaluation, we record the success rate.
Table 11: Total rewards, average speed, and success rate for different $\omega$
@@ 6 Conclusion
Based on knowledge transfer techniques, this paper proposes a novel framework named S2CD. The proposed framework first trains a teacher model within a simplified simulation environment and then utilizes it to guide the student agent in a more complex simulation environment, thereby enhancing the safety and efficiency of the student training process. Subsequently, we introduce an innovative RL algorithm called ACPPO, which leverages samples generated from both teacher and student policies while dynamically adjusting the clipping factor based on sample importance, thereby improving learning efficiency. Moreover, we incorporate the KL divergence between the teacher’s and student’s policies as a constraint in model updates, which we solve using the Lagrangian method, enabling the student agent to quickly assimilate the teacher’s policy. To further enhance the student agent’s performance, we designed a gradual weaning strategy to reduce teacher intervention. This approach ensures that, during the later stages of training, the student agent progressively reduces its reliance on the teacher model and autonomously explores more optimal policies. Experimental results in highway lane-change scenarios demonstrate that, compared to traditional RL, Safe RL, Offline RL, IL, and TSF algorithms, the proposed S2CD framework significantly enhances both learning efficiency and model performance. Additionally, it substantially reduces training costs. More importantly, guidance from the teacher model greatly improves safety during training, even when the teacher model performs suboptimally. Furthermore, the S2CD framework is applicable not only to knowledge transfer across simulation environments of varying complexity but also provides a foundation for research on knowledge transfer between simulation and real-world environments.
All experiments in this study were conducted in simulated environments. Although the simulation platform offers safe, controlled, and efficient testing conditions, it cannot fully replicate the complexity and unpredictability of real-world driving environments. While simulation experiments are valuable during the initial research stages, real-world experiments are necessary to further assess the effectiveness and practicality of the proposed framework. Therefore, future research should incorporate real-world experiments to validate the algorithm’s practical utility under actual driving conditions. Furthermore, extending the research to mixed traffic scenarios involving both autonomous and human-driven vehicles will enhance the algorithm’s applicability and generalization capabilities.

### v3

@@ 3.3.2 High-level Decision-Making Layer
Table 1: Parameters of S2CD framework
Hyper-parameters Value Hyper-parameters Value
Maximal learning rate 0.0005 Discount factor $\gamma$ 0.96
Learning rate decay True Lambda entropy $\beta$ 0.01
Total steps of one episode 5,000 Clip parameter $\epsilon$ 0.2
Total training steps 300K Lambda advantage $\lambda$ 0.98
Optimizer AdamW Hyperparameter $q_{1}$ 3
Mini batch size 64 Hyperparameter $q_{2}$ 10
Hyperparameter of AC $\psi$ 0.2 Initial Lagrange multiplier $\xi$ 0.01
Efficiency weight $\alpha_{1}$ 0.5 Safety weight $\alpha_{2}$ 1.0
Tolerance coefficient $\varepsilon$ 0.5
@@ 5.1 Model Training
Figure 4: The training curves for all algorithms are presented. Each algorithm is trained with three different random seeds, accumulating a total of 500K training steps for each case, except for S2CD, which requires only 300K training steps. To ensure a comprehensive assessment of algorithm performance, two evaluation episodes are conducted for every 5,000 training steps. The average value of these two evaluation episodes is taken as the result.
@@ 5.2 Model Evaluation
Figure 5: The curves of model evaluation for all algorithms. Each model trained with varying random seeds is evaluated twice, and each algorithm undergoes six evaluation runs, with each evaluation consisting of 200 episodes.
@@ 5.3.1 The Impact of Different Teacher Quality
Figure 6: The training and evaluation processes involve different teacher models. 1. Model Training: Each algorithm is trained with three different random seeds, accumulating a total of 300K training steps for each case. Two evaluation episodes are conducted for every 5,000 training steps. The average value of these two evaluation episodes is taken as the result. During model training, we record the return value, safety cost, and collision counts. 2. Model Evaluation: Each model trained with varying random seeds is evaluated twice. Each algorithm undergoes six evaluation runs, with each evaluation consisting of 200 episodes. For model evaluation, we record the success rate.
Table 10: The performance of S2CD under the guidance of teachers at different levels
@@ 5.3.2 Ablations of Guardian Mechanism
Figure 7: The training and evaluation processes of the ablation experiment. 1. Model Training: Each algorithm is trained with three different random seeds, accumulating a total of 300K training steps for each case. Two evaluation episodes are conducted for every 5,000 training steps, and the average value of these two episodes is taken as the result. During model training, we record the return value, safety cost, speed, and collision counts. 2. Model Evaluation: Each model trained with varying random seeds is evaluated twice, and each algorithm undergoes six evaluation runs, with each evaluation consisting of 200 episodes. For model evaluation, we record the success rate.
Table 11: The performance of S2CD in ablation experiments
@@ 6 Conclusion
This paper introduces a novel framework, S2CD, based on knowledge transfer techniques. The S2CD framework first trains a teacher model in a simplified simulation environment and then utilizes this model to guide the student agent in a more complex environment, enhancing the safety and efficiency of the training process. The paper also presents an innovative RL algorithm, ACPPO, which leverages samples generated by both teacher and student policies while dynamically adjusting the clipping factor based on sample importance, thereby improving learning efficiency. Moreover, the KL divergence between the teacher’s and student’s policies is incorporated as a constraint in model updates, solved using the Lagrangian method to enable the student agent to quickly adopt the teacher’s policy. To further enhance the student agent’s performance, a gradual weaning strategy reduces teacher intervention, ensuring that in the later stages of training, the student becomes more autonomous and explores optimal policies independently. Experimental results in highway lane-change scenarios demonstrate that compared to traditional RL, Safe RL, Offline RL, IL, and TSF algorithms, the S2CD framework significantly enhances both learning efficiency and model performance, while also reducing training costs. Most importantly, even with suboptimal teacher performance, the S2CD framework enhances safety during training. The framework is also applicable beyond varying simulation environments and offers a foundation for future knowledge transfer research between simulation and real-world environments.
The primary limitation of this study is that all experiments were conducted in simulated environments, without utilizing teacher models trained in simulation environments to guide student models in real-world conditions. Therefore, the next step in our research will involve real-world vehicle experiments to further validate the effectiveness and practicality of the proposed framework. Additionally, this study did not consider mixed traffic scenarios involving both autonomous and non-autonomous vehicles, which represents another crucial direction for future research to expand the applicability and coverage of the S2CD framework.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC) under grant number 51975194.